# arXiv:2511.10788v2[cs.AI] 01 Mar 2026

### v2

@@ 3.2.3 Modular Methods
4 Conclusion
This survey reframes reasoning in large language models through the lens of adaptivity rather than efficiency alone. It defines adaptivity as the capability to dynamically modulate reasoning depth, structure, or style according to task complexity and uncertainty, provides cognitive formulations of inductive, deductive, and abductive reasoning within LLMs, and introduces a unified taxonomy that distinguishes training-based and training-free implementations. Together, these contributions establish a coherent framework for analyzing and advancing adaptive reasoning toward more efficient, flexible, and context-sensitive language models.

### v3

# From Efficiency to Adaptivity: A Deeper Look at Adaptive Reasoning in Large Language Models

@@ 3.2.3 Modular Methods
4 Conclusion and Open Challenges
This survey reframes reasoning in LLMs through the lens of adaptivity, defining it as the dynamic modulation of reasoning resources based on input-dependent complexity. We formalize this via classical cognitive paradigms and a unified taxonomy of training-based and training-free mechanisms.
However, as the field shifts toward test-time scaling exemplified by models like OpenAI o1 Jaech et al. (2024) and DeepSeek-R1 Guo et al. (2025), new challenges emerge. A primary concern is the reliability of long-chain reasoning; while scaling inference computing can unlock complex problem-solving, it frequently leads to “reasoning drift,” where minor errors accumulate into hallucinations. Future adaptivity mechanisms must therefore transition from simple halting to active error-aware backtracking and dynamic pruning of reasoning paths. Furthermore, the meta-reasoning overhead remains a bottleneck; adaptation signals must be lightweight, potentially utilizing embedding-level difficulty estimations, to ensure that the cost of deciding how to reason does not offset the efficiency gains. Crucially, extending these mechanisms to dynamic and embodied systems represents a major frontier. Recent work such as Ares Yang et al. (2026) and RARRL Liu et al. (2026) demonstrate that agents must autonomously balance high-level cognition with execution constraints, while others adapt via continuous environmental feedback Kuroki et al. . Finally, as adaptive reasoning extends into creative and exploratory domains, the community requires more nuanced metrics that go beyond exact-match accuracy to capture the marginal utility and semantic richness of test-time compute.